Copper(I) oxide can be oxidized to copper(II) oxide: Cu2O(s) + 1⁄2 O2(g) →2CuO(s) ∆H°rxn = - 146.0 kJ

Given that ∆H°f of Cu2O(s) = -168.6 kJ/mol, what is ∆H°f of CuO(s)?


1
Expert's answer
2021-09-26T12:14:24-0400

"Cu_2O(s) +\frac{ 1}{2}O_2(g) \to 2CuO(s)"


"\u2206H\u00barxn = -146.0 kJ"


"∆Hºrxn = 2 ×∆Hf CuO(s) - [\frac{1}{2 }∆Hºf O_
2(g) + ∆Hºf Cu_2O]"


-146 = 2 x ∆Hf CuO(s) - (0 + -168.6)

-146 - 168.6 = 2 x ∆Hf CuO(s)

314.6 kJ = 2 x ∆Hf CuO(s)

∆Hf CuO(s) = 157.3 kJ/mol